OpenStudy (anonymous):

How many solutions does the equation 6s – 3s – 9 = –2 + 3 have? (1 point) Only one None Two Infinitely many

OpenStudy (anonymous):

combine like terms what do you get?

OpenStudy (anonymous):

is it clear what "combine like terms" means?

OpenStudy (anonymous):

one, s = 10/3

OpenStudy (hockeychick23):

6s – 3s – 9 = –2 + 3 3s-9= 1 Add 9 to both sides 3s=10 Divide both sides by 3 s= 10/3